OpenStudy (anonymous):

8/t divided by 16/t?

OpenStudy (amistre64):

multiply top and bottom by t/16

OpenStudy (anonymous):

(8/t)/(16/t) = (8/t)*(t/16) = 1/2

OpenStudy (anonymous):

no t in answer?

OpenStudy (anonymous):

you multiply 8/t by the reciprocal of 16/t, the ts cancel out

OpenStudy (amistre64):

the ts cancel out in the end

OpenStudy (anonymous):

oooookay thanx guys

OpenStudy (anonymous):

\[\frac{8}{t} \div \frac{16}{t}=\frac{8}{t}*\frac{t}{16}\]

OpenStudy (amistre64):

\[\frac{\frac{8}{t}}{\frac{16}{t}}*\frac{\frac{t}{16}}{\frac{t}{16}}=\frac{\frac{8}{16}}{1}\]